A hair dye is made by blending a 8% hydrogen peroxide solution and a 3% hydrogen peroxide solution. How many milliliters of each are used to make a 500-milliliter solution that is 4% hydrogen peroxide?
Answer by checkley77(12844)

.08X+.03(500-X)=.04*500
.08X+15-.03X=20
.05X=20-15
.05X=5
X=5/.05
X=100 ML. OF 8% SOLUTION IS USED.
500-100=400
PROOF:
.08*100+.03(500-100)=.04*500
8+.03*400=20
5+12=20
20=20